# An introduction to infinite projected entangled-pair state methods for variational ground state simulations using automatic differentiation

### v3

@@ Abstract
I Introduction
@@
The content of this work is organised in three main sections. In Sec. II, we describe the central methods that are being used in the variational iPEPS framework as well as practical remarks regarding implementation. Furthermore, we explain in detail the basics of automatic differentiation and its application in state-of-the-art ground-state search. In Sec. III, we then turn to explaining how to conveniently map generic lattice structures to a square one, over which the variational iPEPS methods naturally operate. Following up on this, in Sec. IV, we present numerical benchmarks obtained with the methods outlined in the previous sections and implemented in the variPEPS library, in comparison to other customary methods like exact diagonalization, iPEPS imaginary-time evolution and variational Monte Carlo methods.
II Variational iPEPS
We seek to find the the TN representation of the state vector |$\ket{\psi}_{\mathrm{TN}}$ that best approximates the true ground state vector |$\ket{\psi_{0}}$ of an Hamilton operator of the form
@@ II Variational iPEPS
The method we will present in this section gives rise to an upper bound of the ground state energy in the sense of the variational principle as stated in Eq. (2). But we wish to point out at this point that for that to be strictly true it would be necessary to choose the CTMRG refinement parameter $\chi_{E}$, introduced in detail in Sec. II.2, to be $\chi_{E}\to\infty$. However, in practice we increase this refinement parameter $\chi_{E}$ until all observables are converged.
II.1 iPEPS setup
@@ II.2 CTMRG backbone
In one full CTMRG step, the complete iPEPS unit cell is absorbed into the four lattice directions, such that the eight CTMRG tensors are updated for every iPEPS tensor. This is done column-by-column or row-by-row, depending on the direction. In each absorption step the environment bond dimension $\chi_{E}$ grows by a factor of $\chi_{B}^{2}$. To avoid an exponential increase in memory consumption and computation time, we need a method to truncate the bond dimension back to $\chi_{E}$. In order to do this, we calculate renormalization projectors for each row or column. Projectors are computed from a suitable patch of the iPEPS state including the effective environments, to find a best-possible truncation of the bond dimension. Different approaches for their calculations have been proposed in the literature, which we will discuss in detail below, especially in the context of AD. In the following description of the CTMRG procedure we focus on a left absorption move, which grows all left environment tensors $\{C_{4},T_{4},C_{1}\}$. The main steps of insertion, absorption and renormalization are shown in Fig. 3. In Sec. II.2.1, we will explain the full absorption procedure including renormalization, as it is done in practise. Although projectors need to be calculated before the absorption, their motivation and the calculation of different projects is discussed later in Sec. II.2.2.
II.2.1 Absorption of iPEPS tensors
@@
The absorption of a full unit cell is then performed for the other three directions. In a top move the tensors $C_{1}$, $T_{1}$ and $C_{2}$ are grown, in a right move the tensors $C_{2}$, $T_{2}$ and $C_{3}$ and in a bottom move the tensors $C_{3}$, $T_{3}$ and $C_{4}$. This completes a single CTMRG step, which is then repeated in the directional procedure until convergence is reached. In Sec. II.2.3 we discuss appropriate convergence measures.
II.2.2 Calculation of projectors
@@
II.2.3 Convergence and CTMRG fixed-points
@@
One peculiarity that is however not incorporated in this convergence check is sign or phase fluctuation for real or complex tensor entries, respectively. This means that, while projectors and hence the CTMRG tensors converge in absolute value, their entries can have different signs/phases in consecutive CTM steps. For reasons that become clear in Sec. II.5 it is however required to reach element-wise convergence in the environment tensors for them to represent an actual fixed-point [42]. Those fluctuations originate from the gauge freedom in the SVD performed in Eq. (6). This is reflected in the freedom of introducing a unitary (block-)diagonal matrix $\Gamma$ in an SVD,
@@
II.3 Energy expectation values
@@ II.4 Automatic differentiation
In the forward-mode AD scheme, the gradient is built up simultaneously with the evaluation of the primitives $f_{1},f_{2}$ and $f_{3}$, according to the following prescription for the $i$-th step:
$\begin{split}f_{i}(\vec{x}_{i}^{0})&=\vec{x}_{i+1}^{0}\\ G_{i}&=J^{i}(\vec{x}_{i}^{0})\cdot G_{i-1}\end{split}$ (16)
with the starting condition $G_{0}:=\mathds{1}_{n_{1}\times n_{1}}$ and the final result $G_{3}=\nabla E(\vec{x}_{1}^{0})\in\mathbb{R}^{n_{4}\times n_{1}}$. We see that in this case we build up Eq. (15) from right to left or “along the computational graph” as illustrated in Fig. 13.
@@
II.5 Calculation of the gradient at the CTMRG fixed-point
Computationally, the CTMRG routine represents the bottleneck of the full iPEPS energy function. It involves many expensive contractions and SVDs. Moreover, it requires an a priori unknown number of CTMRG iterations to reach convergence of the environment tensors. This would be especially disadvantageous for the gradient evaluation using plain-vanilla backward-mode AD, since this would require unrolling all the performed CTMRG iterations and paying a memory consumption linear in their number. However, this can be avoided by leveraging that fact that the CTMRG iteration eventually converges to a fixed point, and this is precisely the condition under which the energy evaluation is then performed. As soon this fixed point is reached, all CTMRG iterations are identical, i.e., reproducing the converged environment tensors. We can, in this situation, get away with only saving intermediate results from such a converged CTMRG iteration. This reduces the memory requirements by a factor of the number of CTMRG iterations that we perform [37]. We stress here that, for this approach to work, we must make sure that the CTMRG procedure reaches an actual fixed point, meaning that all CTMRG environment tensors are converged element wise as discussed in Sec. II.2.3. The fixed-point equation can be written as
@@ II.5 Calculation of the gradient at the CTMRG fixed-point
II.6 Optimization
As discussed in the introduction of Sec. II we seek to find the iPEPS approximation |$\ket{\psi}_{\text{TN}}$ of the ground state vector |$\ket{\psi_{0}}$. Employing the methods discussed in the last sections we can describe this energy calculation as function $E(\ket{\psi}_{\text{TN}})$, consisting of the CTMRG power-method and the expectation value approximation using the resulting CTMRG environment tensors. Since we can calculate the gradient $\nabla E(\ket{\psi}_{\text{TN}})$ of this real scalar function it is straightforward to use well-known optimization methods to find the energy minimum. We would like to stress that the state vector |$\ket{\psi}_{\text{TN}}$, and thus the energy function, only depends on the tensors defining the iPEPS ansatz and not the environment tensors since they are implicitly calculated from the ansatz. In this discussion we focus on two types of methods based on the gradient: The (nonlinear) conjugate gradient (CG) [66, 67, 68, 69, 70] and the quasi-Newton methods [71, 72, 73, 74, 75, 76].
@@ II.6 Optimization
II.7 Pitfalls and practical hints
II.7.1 Iterative SVD algorithm
We also advertise the use of iterative algorithms for the calculation of the SVD in the CTMRG procedure. This can be quite advantageous computationally, since only $\chi_{E}$ singular values are needed for a matrix of size $(\chi_{E}\chi_{B}^{2})\times(\chi_{E}\chi_{B}^{2})$ during the CTMRG. To this end we use the use the Golub-Kahan-Lanczos (GKL) bidiagonalization algorithm with additional orthogonalization for the Krylov vectors. This algorithm is available, e.g., in packages like KrylovKit.jl [83] or IterativeSolvers.jl [84] in the Julia programming language. We highlight the utility of this type of algorithm for the calculation of the SVD with the comparison of the computational time of the different algorithms in the iPEPS use case in Fig. 16.
@@ II.7.1 Iterative SVD algorithm
Figure 16: Comparison of the computational time for the calculation of the first $\chi_{E}$ singular values/vectors of a matrix of dimension $(\chi_{E}\chi_{B}^{2})\times(\chi_{E}\chi_{B}^{2})$ obtained in a CTMRG procedure with bond dimension $\chi_{B}=6$. The conventional SVD (blue), which is truncated only after calculating the full SVD spectrum is substantially slower than the iterative GKL methods. The GKL algorithm in the CTMRG use case was showed comparable performance when constructing the $\chi_{E}\chi_{B}^{2}$ matrix explicitly (orange) or by just implementing its action of a vector (green). While the GKL algorithm for the case at moderate $d$ and $\chi_{E}$ constructing the matrix usually is faster, at larger $\chi_{B}$ and $\chi_{E}$ it can become advantages to only implement the action of the matrix.
II.7.2 Stability of the CTMRG routine
One of the basic prerequisite for a stable variational iPEPS optimization is a robust CTMRG routine fulfilling the convergence requirements discussed in Sec. II.2.3. Obviously, there is the environment bond dimension $\chi_{E}$ to control the accuracy of the approximation of the environment. If the environment bond dimension is chosen too low, the approximation is invalid and the CTMRG routine can yield an inaccurate result for the expectation value. This could further lead to an unstable variational update. To check heuristically whether the refinement parameter $\chi_{E}$ is chosen sufficiently high, one can check the singular value spectrum obtained during the projector calculation as described in Sec. II.2.2. As a reliable criteria for the amount of information loss, we compute the truncation error $\varepsilon_{T}$ given by the norm of the discarded singular values of the normalized spectrum [85]. If the truncation error is larger than some threshold (e.g., $\varepsilon_{T}>10^{-5}$), one can assume that the environment bond dimension is chosen too low and has to be increased. Employing this procedure, the bond dimension can automatically be increased during the variational optimization if necessary. A sufficiently large $\chi_{E}$ is crutial as the AD optimization can otherwise exploit the inaccuracies of the CTMRG procedure, leading to false ground states with artificially low energy.
II.7.3 Prevention of local minima
@@
An optimization of a system with a tendency for local minima might still be successful, if a suitable initial state is provided. One possibility are initial states obtained by imaginary-time evolution methods (simple update, full update [86, 22, 23]). While this is typically a convenient solution, it is sometimes necessary to perturb the input tensors with a small amount of noise (e.g., $10^{-2}$ in relative amplitude) to actually avoid local minima. As an alternative, one can input a converged state obtained from energy minimization of a different TN ansatz, provided there is a suitable mapping between the different structures. Examples for this technique are provided for benchmarks on different lattices in Sec. IV.
@@ II.7.4 Recycling of environments
II.7.5 Analysing iPEPS data at finite bond dimensions
@@
II.7.6 Degenerate singular values
@@
III Extension to other lattices
The directional CTMRG routine on the square lattice is very convenient for its orthogonal lattice vectors and definition of the effective environments. It is therefore natural to exploit the implemented routines for different kind of lattices that can be mapped back to the square lattice. This can typically be achieved by a suitable coarse-graining, in which a collection of lattice sites on the original lattice is mapped into an effective site on the square lattice. Energy expectation values can then be directly evaluated in the coarse-grained picture as well. This is even advantageous for the AD optimization procedure, since the energy can often be computed with a smaller number of individual terms. In this section we will present the mapping for four types of lattices frequently found in condensed matter systems – the honeycomb, Kagome, square-Kagome and triangular lattice. Naturally, the framework can be extended by other suitable two-dimensional lattices, such as dice, square-octagon, maple-leaf and others. As an alternative to the coarse-graining approach, CTMRG methods that directly operate on the original lattice structures can also be defined [46, 93].
III.1 Honeycomb lattice

### v4

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics Lecture Notes.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
1 Introduction
@@
The content of this work is organised in three main sections. In Sec. 2, we describe the central methods that are being used in the variational iPEPS framework as well as practical remarks regarding implementation. Furthermore, we explain in detail the basics of automatic differentiation and its application in state-of-the-art ground-state search. In Sec. 3, we then turn to explaining how to conveniently map generic lattice structures to a square one, over which the variational iPEPS methods naturally operate. Following up on this, in Sec. 4, we present numerical benchmarks obtained with the methods outlined in the previous sections and implemented in the variPEPS library, in comparison to other customary methods like exact diagonalization, iPEPS imaginary-time evolution and variational Monte Carlo methods.
2 Variational iPEPS
We seek to find the TN representation of the state vector |$\ket{\psi}_{\mathrm{TN}}$ that best approximates the true ground state vector |$\ket{\psi_{0}}$ of an Hamilton operator of the form
@@ 2 Variational iPEPS
Besides the CTMRG procedure, other well-controlled numerical methods have been developed to contract the infinite PEPS network. These include boundary matrix product states [22, 59] and tensor coarse-graining techniques [27, 28, 29, 30]. These methods can be equally successfully combined with the concept of automatic differentiation [37, 47] and provide competitive results. While other methods can be generalized to non-trivial unit cells [60] as well, in this work we focus on the CTMRG method due to its straightforward generalization and flexibility in handling arbitrary unit cells of size $(L_{x},L_{y})$. Furthermore, the CTMRG method provides a robust agnosticism towards possible non-Hermitian transfer operators, which has to be treated with more care in, e.g., boundary MPS methods.
The method we will present in this section gives rise to an upper bound of the ground state energy in the sense of the variational principle as stated in Eq. (2). But we wish to point out at this point that for that to be strictly true it would be necessary to choose the CTMRG refinement parameter $\chi_{E}$, introduced in detail in Sec. 2.2, to be $\chi_{E}\to\infty$. However, in practice we increase this refinement parameter $\chi_{E}$ until all observables are converged.
2.1 iPEPS setup
@@ 2.2 CTMRG backbone
In one full CTMRG step, the complete iPEPS unit cell is absorbed into the four lattice directions, such that the eight CTMRG tensors are updated for every iPEPS tensor. This is done column-by-column or row-by-row, depending on the direction. In each absorption step the environment bond dimension $\chi_{E}$ grows by a factor of $\chi_{B}^{2}$. To avoid an exponential increase in memory consumption and computation time, we need a method to truncate the bond dimension back to $\chi_{E}$. In order to do this, we calculate renormalization projectors for each row or column. Projectors are computed from a suitable patch of the iPEPS state including the effective environments, to find a best-possible truncation of the bond dimension. Different approaches for their calculations have been proposed in the literature, which we will discuss in detail below, especially in the context of AD. In the following description of the CTMRG procedure we focus on a left absorption move, which grows all left environment tensors $\{C_{4},T_{4},C_{1}\}$. The main steps of insertion, absorption and renormalization are shown in Fig. 3. In Sec. 2.2.1, we will explain the full absorption procedure including renormalization, as it is done in practise. Although projectors need to be calculated before the absorption, their motivation and the calculation of different projects is discussed later in Sec. 2.2.2.
2.2.1 Absorption of iPEPS tensors
@@
The absorption of a full unit cell is then performed for the other three directions. In a top move the tensors $C_{1}$, $T_{1}$ and $C_{2}$ are grown, in a right move the tensors $C_{2}$, $T_{2}$ and $C_{3}$ and in a bottom move the tensors $C_{3}$, $T_{3}$ and $C_{4}$. This completes a single CTMRG step, which is then repeated in the directional procedure until convergence is reached. In Sec. 2.2.3 we discuss appropriate convergence measures.
2.2.2 Calculation of projectors
@@
2.2.3 Convergence and CTMRG fixed-points
@@
One peculiarity that is however not incorporated in this convergence check is sign or phase fluctuation for real or complex tensor entries, respectively. This means that, while projectors and hence the CTMRG tensors converge in absolute value, their entries can have different signs/phases in consecutive CTM steps. For reasons that become clear in Sec. 2.5 it is however required to reach element-wise convergence in the environment tensors for them to represent an actual fixed-point [42]. Those fluctuations originate from the gauge freedom in the SVD performed in Eq. (6). This is reflected in the freedom of introducing a unitary (block-)diagonal matrix $\Gamma$ in an SVD,
@@
2.3 Energy expectation values
@@ 2.4 Automatic differentiation
In the forward-mode AD scheme, the gradient is built up simultaneously with the evaluation of the primitives $f_{1},f_{2}$ and $f_{3}$, according to the prescription
$\begin{split}f_{i}(\vec{x}_{i}^{0})&=\vec{x}_{i+1}^{0},\\ G_{i}&=J^{i}(\vec{x}_{i}^{0})\cdot G_{i-1}\end{split}$ (16)
for the $i$-th step, with the starting condition $G_{0}:=\mathds{1}_{n_{1}\times n_{1}}$ and with the final result that is given by $G_{3}=\nabla E(\vec{x}_{1}^{0})\in\mathbb{R}^{n_{4}\times n_{1}}$. We see that in this case we build up Eq. (15) from right to left or “along the computational graph” as illustrated in Fig. 13.
@@
2.5 Calculation of the gradient at the CTMRG fixed-point
Computationally, the CTMRG routine represents the bottleneck of the full iPEPS energy function. It involves many expensive contractions and SVDs. Moreover, it requires an a priori unknown number of CTMRG iterations to reach convergence of the environment tensors. This would be especially disadvantageous for the gradient evaluation using plain-vanilla backward-mode AD, since this would require unrolling all the performed CTMRG iterations and paying a memory consumption linear in their number. However, this can be avoided by leveraging that fact that the CTMRG iteration eventually converges to a fixed point, and this is precisely the condition under which the energy evaluation is then performed. As soon this fixed point is reached, all CTMRG iterations are identical, i.e., reproducing the converged environment tensors. We can, in this situation, get away with only saving intermediate results from such a converged CTMRG iteration. This reduces the memory requirements by a factor of the number of CTMRG iterations that we perform [37]. We stress here that, for this approach to work, we must make sure that the CTMRG procedure reaches an actual fixed point, meaning that all CTMRG environment tensors are converged element wise as discussed in Sec. 2.2.3. The fixed-point equation can be written as
@@ 2.5 Calculation of the gradient at the CTMRG fixed-point
2.6 Optimization
As discussed in the introduction of Sec. 2 we seek to find the iPEPS approximation |$\ket{\psi}_{\text{TN}}$ of the ground state vector |$\ket{\psi_{0}}$. Employing the methods discussed in the last sections we can describe this energy calculation as function $E(\ket{\psi}_{\text{TN}})$, consisting of the CTMRG power-method and the expectation value approximation using the resulting CTMRG environment tensors. Since we can calculate the gradient $\nabla E(\ket{\psi}_{\text{TN}})$ of this real scalar function it is straightforward to use well-known optimization methods to find the energy minimum. We would like to stress that the state vector |$\ket{\psi}_{\text{TN}}$, and thus the energy function, only depends on the tensors defining the iPEPS ansatz and not the environment tensors since they are implicitly calculated from the ansatz. In this discussion we focus on two types of methods based on the gradient: The (nonlinear) conjugate gradient (CG) [68, 69, 70, 71, 72] and the quasi-Newton methods [73, 74, 75, 76, 77, 78].
@@ 2.6 Optimization
2.7 Handling of physical symmetries
One important property of tensor networks is their ability to incorporate physical symmetries, such as global internal symmetries or spatial symmetries, into their structure exactly. This can be achieved at the level of each individual tensor, ensuring that the entire many-body wave function remains symmetry-invariant. On one hand, this approach allows for the targeting of specific sectors of the Hamiltonian and facilitates a symmetry-resolved analysis of the model under consideration. On the other hand, it restricts the number of remaining variational parameters and can lead to a substantial computational speed-up, which, in turn, enables access to larger bond dimensions.
Global symmetries can be implemented by making the tensors quantum number-preserving, with quantum numbers corresponding to the underlying symmetry group. This results in a sparse tensor block structure, where linear algebra operations are performed on a generally larger set of much smaller individual tensors, leading to more efficient manipulations. Common symmetries exploited in tensor network simulations include Abelian symmetries such as $\mathbb{Z}_{N}$ or $U(1)$ [85, 86], non-Abelian symmetries such as $SU(2)$ [87, 88], and also fermionic symmetries [89, 90]. Moreover, lattice symmetries can be directly implemented as well, e.g. by imposing reflection or rotation symmetry for the tensor entries.
By making the full set of algorithmic tensor operations, such as tensor initialization, contraction, factorization, index permutation, etc., aware of the block structure, symmetric tensors can be treated analogously to non-symmetric ones on a conceptual level. This includes their use in variational optimization based on automatic differentiation. There exist several open source libraries that can handle internal (non)-Abelian and lattice symmetries in TNs [91, 92, 93, 94], some already within the scope of variational PEPS optimization [50].
2.8 Pitfalls and practical hints
2.8.1 Iterative SVD algorithm
We also advertise the use of iterative algorithms for the calculation of the SVD in the CTMRG procedure. This can be quite advantageous computationally, since only $\chi_{E}$ singular values are needed for a matrix of size $(\chi_{E}\chi_{B}^{2})\times(\chi_{E}\chi_{B}^{2})$ during the CTMRG. To this end, we use the Golub-Kahan-Lanczos (GKL) bidiagonalization algorithm with additional orthogonalization for the Krylov vectors. This algorithm is available, e.g., in packages like KrylovKit.jl [95] or IterativeSolvers.jl [96] in the Julia programming language. We highlight the utility of this type of algorithm for the calculation of the SVD with the comparison of the computational time of the different algorithms in the iPEPS use case in Fig. 16.
@@ 2.8.1 Iterative SVD algorithm
Figure 16: Comparison of the computational time for the calculation of the first $\chi_{E}$ singular values/vectors of a matrix of dimension $(\chi_{E}\chi_{B}^{2})\times(\chi_{E}\chi_{B}^{2})$ obtained in a CTMRG procedure with bond dimension $\chi_{B}=6$. The conventional SVD (blue), which is truncated only after calculating the full SVD spectrum is substantially slower than the iterative GKL methods. The GKL algorithm in the CTMRG use case was showed comparable performance when constructing the $\chi_{E}\chi_{B}^{2}$ matrix explicitly (orange) or by just implementing its action of a vector (green). While the GKL algorithm for the case at moderate $d$ and $\chi_{E}$ constructing the matrix usually is faster, at larger $\chi_{B}$ and $\chi_{E}$ it can become advantageous to only implement the action of the matrix.
2.8.2 Stability of the CTMRG routine
One of the basic prerequisite for a stable variational iPEPS optimization is a robust CTMRG routine fulfilling the convergence requirements discussed in Sec. 2.2.3. Obviously, there is the environment bond dimension $\chi_{E}$ to control the accuracy of the approximation of the environment. If the environment bond dimension is chosen too low, the approximation is invalid and the CTMRG routine can yield an inaccurate result for the expectation value. This could further lead to an unstable variational update. To check heuristically whether the refinement parameter $\chi_{E}$ is chosen sufficiently high, one can check the singular value spectrum obtained during the projector calculation as described in Sec. 2.2.2. As a reliable criteria for the amount of information loss, we compute the truncation error $\varepsilon_{T}$ given by the norm of the discarded singular values of the normalized spectrum [97]. If the truncation error is larger than some threshold (e.g., $\varepsilon_{T}>10^{-5}$), one can assume that the environment bond dimension is chosen too low and has to be increased. Employing this procedure, the bond dimension can automatically be increased during the variational optimization if necessary. A sufficiently large $\chi_{E}$ is crucial as the AD optimization can otherwise exploit the inaccuracies of the CTMRG procedure, leading to false ground states with artificially low energy.
2.8.3 Prevention of local minima
@@
An optimization of a system with a tendency for local minima might still be successful, if a suitable initial state is provided. One possibility are initial states obtained by imaginary-time evolution methods (simple update, full update [98, 22, 23]). While this is typically a convenient solution, it is sometimes necessary to perturb the input tensors with a small amount of noise (e.g., $10^{-2}$ in relative amplitude) to actually avoid local minima. As an alternative, one can input a converged state obtained from energy minimization of a different TN ansatz, provided there is a suitable mapping between the different structures. Examples for this technique are provided for benchmarks on different lattices in Sec. 4.
@@ 2.8.4 Recycling of environments
2.8.5 Analysing iPEPS data at finite bond dimensions
@@
2.8.6 Degenerate singular values
@@
3 Extension to other lattices
The directional CTMRG routine on the square lattice is very convenient for its orthogonal lattice vectors and definition of the effective environments. It is therefore natural to exploit the implemented routines for different kind of lattices that can be mapped back to the square lattice. This can typically be achieved by a suitable coarse-graining, in which a collection of lattice sites on the original lattice is mapped into an effective site on the square lattice. Energy expectation values can then be directly evaluated in the coarse-grained picture as well. This is even advantageous for the AD optimization procedure, since the energy can often be computed with a smaller number of individual terms. In this section we will present the mapping for four types of lattices frequently found in condensed matter systems – the honeycomb, Kagome, square-Kagome and triangular lattice. Naturally, the framework can be extended by other suitable two-dimensional lattices, such as dice, square-octagon, maple-leaf and others. As an alternative to the coarse-graining approach, CTMRG methods that directly operate on the original lattice structures can also be defined [46, 105, 106].
3.1 Honeycomb lattice